# High-resolution ultra-low-field MRI with SNRAware denoising

### 02 Apr 2026 v1

@@ IV-G Model performance versus anatomy, orientation, and contrast
The results of Experiment 7 demonstrate that SNRAware provides substantial SNR improvements across a broad range of anatomical targets, imaging orientations, and contrasts. As shown in Fig. 6, the denoised reconstructions exhibit visibly higher SNR than the corresponding raw FFT images for both knee and wrist acquisitions, and across T1-w, T2-w, and STIR contrasts. Similar behavior has also been reported in ULF brain imaging [31]. The improvement is consistent across axial, sagittal, and coronal orientations, indicating that the model generalizes well to variations in anatomical structure and image geometry. The amount of detail recovered in the STIR denoised image ($\text{TI}=80$ ms) is particularly remarkable.
@@ IV-H Model performance versus BM4D
From the results of Experiment 8 it is apparent that classical BM4D filtering also provides a strong improvement in terms of SNR in the selected ROIs, with boosts ranging from 3.1 to 26.9 ($8.7\times$) for T1-w, to 2.5 to 26.2 ($10.5\times$) for STIR. These are to be compared with SNR boosts of $9.4\times$ and $15.6\times$ in the corresponding DL-denoised reconstructions. However, this is for the case where $\sigma_{\text{BM4D}}=2$, which leads to significant blurring and loss of anatomical detail (see red arrows). These effects are less evident with $\sigma_{\text{BM4D}}=1$, but the denoising performance is significantly lower, with SNR increases of only $1.9\times$ and $5.9\times$ for the T1-w and STIR images, respectively. These results show strong advantages of SNRAware with respec to BM4D in low-SNR imaging regimes.
@@ V Conclusions & Outlook
A central observation emerging from this study is the ability of the model to remove noise while preserving the underlying signal structure. Residual images obtained by subtracting the denoised reconstructions from the raw images typically exhibit noise-like patterns with minimal structured content, indicating that the model acts predominantly as a noise suppressor rather than performing implicit image restoration or hallucination. Given the fundamentally stochastic nature of noise, the ability of an algorithm to identify and remove noise contributions in this manner may appear counterintuitive. The results presented here nevertheless demonstrate that DL-based models trained with appropriate noise statistics can perform this task with remarkable effectiveness.
At the same time, the experiments highlight several limitations. The quality of the denoised output remains strongly dependent on the starting SNR of the acquisition, as illustrated by the elevated-noise experiment. Likewise, the performance of the model appears to improve as the spatial resolution approaches the regime of the training data, suggesting that training-domain mismatch plays a role in determining the model’s behavior. Hardware factors such as impedance matching, bandwidth selection, and environmental noise also influence the final reconstruction quality, emphasizing that denoising performance ultimately reflects the combined properties of both the acquisition system and the post-processing algorithm.

### 05 Apr 2026 v2

@@ IV-G Model performance versus anatomy, orientation, and contrast
The results of Experiment 7 demonstrate that SNRAware provides substantial SNR improvements across a broad range of anatomical targets, imaging orientations, and contrasts. As shown in Fig. 6, the denoised reconstructions exhibit visibly higher SNR than the corresponding raw FFT images for both knee and wrist acquisitions, and across T1-w, T2-w, and STIR contrasts. Similar behavior has also been reported in ULF brain imaging [31]. The improvement is consistent across axial, sagittal, and coronal orientations, indicating that the model generalizes well to variations in anatomical structure and image geometry. The amount of detail recovered in the STIR image ($\text{TI}=80$ ms) denoised with the model is particularly remarkable. The BM4D denoised images also recover some of the bone structure, but to a lesser extent.
@@ IV-H Model performance versus BM4D
From the results of Experiment 8 it is apparent that classical BM4D filtering also provides a strong improvement in terms of SNR in the selected ROIs, with boosts ranging from 3.1 to 26.9 ($8.7\times$) for T1-w, to 2.5 to 26.2 ($10.5\times$) for STIR (see Fig. 7). These are to be compared with SNR boosts of $9.4\times$ and $15.6\times$ in the corresponding DL-denoised reconstructions. However, this is for the case where $\sigma_{\text{BM4D}}=2$, which leads to significant blurring and loss of anatomical detail (see red arrows). These effects are less evident with $\sigma_{\text{BM4D}}=1$, but the denoising performance is significantly lower, with SNR increases of only $1.9\times$ and $5.9\times$ for the T1-w and STIR images, respectively. These results show strong advantages of SNRAware with respec to BM4D in low-SNR imaging regimes.
@@ V Conclusions & Outlook
A central observation emerging from this study is the ability of the model to remove noise while preserving the underlying signal structure. Residual images obtained by subtracting the denoised reconstructions from the raw images typically exhibit noise-like patterns with minimal structured content, indicating that the model acts predominantly as a noise suppressor rather than performing implicit image restoration or hallucination. Given the fundamentally stochastic nature of noise, the ability of an algorithm to identify and remove noise contributions in this manner may appear counterintuitive. The results presented here nevertheless demonstrate that DL-based models trained with appropriate noise statistics can perform this task with remarkable effectiveness. This performance likely stems from the SNRAware training strategy, which leverages temporal redundancy in cardiac cine data to distinguish signal components that are consistent across frames from noise that is uncorrelated. When applied to 3D ULF acquisitions, a similar separation appears to take place across slices, where spatially coherent structures are preserved while uncorrelated components are efficiently suppressed, even for extremely low-SNR acquisitions (see Fig. 7(b)).
At the same time, the experiments highlight several limitations. The quality of the denoised output remains dependent on the starting SNR of the acquisition, as illustrated by the elevated-noise experiment (Fig. 5(b)). Likewise, the performance of the model appears to improve as the spatial resolution approaches the regime of the training data, suggesting that training-domain mismatch plays a role in determining the model’s behavior (Fig. 4). Hardware factors such as impedance matching, bandwidth selection, and environmental noise also influence the final reconstruction quality, emphasizing that denoising performance ultimately reflects the combined properties of both the acquisition system and the post-processing algorithm.